# Pretraining Decision Transformers with Reward Prediction for In-Context Multi-task Structured Bandit Learning

### v1

@@ 5 Empirical Study: New Actions
Invariant actions: We denote the set of actions fixed across the different tasks in the pretraining in-context dataset as $\mathcal{A}^{\text{inv}}$. Therefore these action features $\mathbf{x}(a)\in\mathbb{R}^{d}$ for $a\in\mathcal{A}^{\text{inv}}$ are fixed across the different tasks $m$. Note that these invariant actions help the transformer $\mathrm{TF}_{\mathbf{w}}$ to learn the latent structure and the reward correlation across the different tasks.
@@
New actions: However, we also want to test how robust is PreDeToR (-$\tau$) to new actions not seen during training time. To this effect, for each task $m\in[M_{\text{pre}}]$ and $m\in[M_{\text{test}}]$ we introduce $A-|\mathcal{A}^{\text{inv}}|$ new actions. For each of these new actions $a\in[A-|\mathcal{A}^{\text{inv}}|]$ we choose the features $\mathbf{x}(m,a)$ randomly from $\mathcal{X}\subseteq\mathbb{R}^{d}$. Note the transformer now trains on a dataset $\mathcal{H}_{m}\subseteq\mathcal{D}_{\text{pre}}\neq\mathcal{D}_{\text{test}}$.
(a) Linear $1$ new action
(b) Non-linear $1$ new action
(c) Linear $5$ new actions
(d) Non-linear $5$ new actions
@@
Figure 2: New action experiments. The horizontal axis is the number of rounds. Confidence bars show one standard error.
@@
(1) The PreDeToR (-$\tau$) learns to exploit the underlying latent structure and reward correlation between the different tasks even when $\mathcal{D}_{\text{test}}\neq\mathcal{D}_{\text{pre}}$ and the number of changing actions per task is small. So PreDeToR (-$\tau$) is more robust to changing actions compared to AD and DPT-greedy.
(2) PreDeToR (-$\tau$) has lower regret than weak demonstrator even when it is trained on $\mathcal{D}_{\text{pre}}\neq\mathcal{D}_{\text{test}}$.
(3) The PreDeToR (-$\tau$) cumulative regret increases as the number of invariant actions decreases. This shows that invariant actions are important for learning the underlying latent structure and reward correlation across tasks.
@@
Experimental Result: We observe these outcomes in Figure 2. We consider the linear bandit setting of horizon $n=50$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=20$, and $d=5$. Here during data collection and during collecting the test data, we randomly select one new action from $\mathbb{R}^{d}$ for each task $m$. So the number of invariant actions is $|\mathcal{A}^{\text{inv}}|=19$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, and AD. The weak performance of DPT-greedy can be attributed to low data per task and inability to adapt when $\mathcal{D}_{\text{pre}}\neq\mathcal{D}_{\text{test}}$. Observe that PreDeToR (-$\tau$) almost matches LinUCB and has lower regret than Thomp. This shows that PreDeToR (-$\tau$) is able to exploit the latent linear structure of the underlying tasks. In Figure 2(b) we show the non-linear bandit setting for horizon $n=50$, $M_{\text{pre}}=100000$, $M_{\text{test}}=200$,$A=6$, $d=2$, and $|\mathcal{A}^{\text{inv}}|=5$. The demonstrator $\pi^{w}$ is the Thomp algorithm. Again, we observe that PreDeToR has lower cumulative regret than DPT-greedy and AD and furthermore improves upon LinUCB which fails to perform well in this non-linear setting due to its assumption of linear rewards. PreDeToR-$\tau$ performs slightly better than PreDeToR in $1$ new arm setting as it conducts more exploration.
In the following two experiments, we test the limit of robustness of PreDeToR (-$\tau$) by decreasing the number of invariant actions. We use $|\mathcal{A}^{\text{inv}}|=15$ for linear, and $|\mathcal{A}^{\text{inv}}|=2$ for the non-linear setting. All the other task parameters are the same as in the previous two experiments. The corresponding results are shown in Figure 2(c) and Figure 2(d). Observe that in the linear setting (where LinUCB is effective), the performance of PreDeToR degrades compared to the case when $|\mathcal{A}^{\text{inv}}|=15$. This shows that as the number of invariant actions decreases, the ability of PreDeToR (-$\tau$) to estimate the underlying latent structure and reward correlation also decreases. In Section A.12, we analyze the exploration of PreDeToR (-$\tau$) in this setting, which shows that the shared latent structure is essential for minimizing cumulative regret.
@@
We also empirically study the test performance of PreDeToR (-$\tau$) in other bandit settings such as bilinear bandits (Section A.3), latent bandits (Section A.4), draw a connection between PreDeToR and Bayesian estimators (Section A.5), and perform sensitivity and ablation studies in Section A.6, A.8, A.9, A.10. We discuss data collection algorithms in Section A.13 and the offline setting in Section A.15. Due to space constraints, we refer the interested reader to the relevant section in the appendices.
6 Theoretical Analysis
@@
In this section, we present a theoretical analysis of how PreDeToR-$\tau$ generalizes to an unknown target task given a set of source tasks. We also show that in low-data regimes the PreDeToR-$\tau$ has a low expected excess risk for the unknown target task as the number of source tasks increases. We proceed as follows: Suppose we have the training data set $\mathcal{H}_{\mathrm{all}}=\{\mathcal{H}_{m}\}_{m=1}^{M_{\text{pre}}},$ where the task $m\sim\mathcal{T}$ with a distribution $\mathcal{T}$ and the task data $\mathcal{H}_{m}$ is generated from a distribution $\mathcal{D}_{\text{pre}}(\cdot|m).$ For illustration purposes, here we consider the training data distribution $\mathcal{D}_{\text{pre}}(\cdot|m)$ where the actions are sampled following soft-LinUCB (Chu et al., 2011). Note that soft-LinUCB produces a distribution over actions (see Section A.14 for a detailed description). Given the loss function in (3), we can define the task $m$ training loss of PreDeToR-$\tau$ as ^$\widehat{\mathcal{L}}_{m}(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}})=\frac{1}{n}\sum_{t=1}^{n}\ell(r_{m,t},\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\widehat{r}_{m,t}(I_{m,t})|\mathcal{H}^{t}_{m}))=\frac{1}{n}\sum_{t=1}^{n}(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\widehat{r}_{m,t}(I_{m,t})|\mathcal{H}^{t}_{m})-r_{m,t})^{2}$. We drop the notation $\mathbf{\Theta},\mathbf{r}$ from $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}$ for simplicity and let $M=M_{\text{pre}}$. We define
@@
Note that in in-context learning, a training sample $\left(I_{t},r_{t}\right)$ impacts all future decisions of the algorithm from time step $t+1$ to $n$. Therefore, we need to control the stability of the input perturbation of the learning algorithm learned by the transformer. We introduce the following stability condition.
@@
(Error stability (Bousquet and Elisseeff, 2002; Li et al., 2023)). Let $\mathcal{H}=\left(I_{t},r_{t}\right)_{t=1}^{n}$ be a sequence in $[A]\times[0,1]$ with $n\geq 1$ and $\mathcal{H}^{\prime}$ be the sequence where the $t^{\prime}$th sample of $\mathcal{H}$ is replaced by $\left(I_{t}^{\prime},r_{t}^{\prime}\right)$. Error stability holds for a distribution $(I,r)\sim\mathcal{D}$ if there exists a $K>0$ such that for any $\mathcal{H},\left(I_{t}^{\prime},r_{t}^{\prime}\right)\in([A]\times[0,1]),t\leq n$, and $\mathrm{TF}\in\mathrm{Alg}$, we have
$\big{\lvert}\mathbb{E}_{(I,r)}\left[\ell(r,\mathrm{TF}(\widehat{r}(I)|\mathcal{H}))-\ell\left(r,\mathrm{TF}(\widehat{r}(I)|\mathcal{H}^{\prime})\right)\right]\big{\rvert}\,\leq\tfrac{K}{n}$
@@
(PreDeToR risk) Suppose error stability 6.1 holds and assume loss function $\ell(\cdot,\cdot)$ is $C$-Lipschitz for all $r_{t}\in[0,B]$ and horizon $n\geq 1$. Let $\widehat{\mathrm{TF}}$ be the empirical solution of (ERM) and $\mathcal{N}(\mathcal{A},\rho,\epsilon)$ be the covering number of the algorithm space $\mathrm{Alg}$ following Definition C.2 and C.3. Then with probability at least $1-2\delta$, the excess MTL risk of PreDeToR-$\tau$ is bounded by $\mathcal{R}_{\mathrm{MTL}}(\widehat{\mathrm{TF}})\leq 4\tfrac{C}{\sqrt{nM}}+2(B+K\log n)\sqrt{\tfrac{\log(\mathcal{N}(\mathrm{Alg},\rho,\varepsilon)/\delta)}{cnM}},$ where $\mathcal{N}(\mathrm{Alg},\rho,\varepsilon)$ is the covering number of transformer $\widehat{\mathrm{TF}}$ and $\epsilon=1/\sqrt{nM}$.
The proof of this theorem is provided in Section C.1. From Theorem 6.2 we see that in low-data regime with a small horizon $n$, as the number of tasks $M$ increases the MTL risk decreases. We further discuss the stability factor $K$ and covering number $\mathcal{N}(\mathrm{Alg},\rho,\varepsilon)$ in Remark C.4, and C.5.
@@ 7 Conclusions, Limitations and Future Works
In this paper, we studied the supervised pretraining of decision transformers in the multi-task structured bandit setting when the knowledge of the optimal action is unavailable. Moreover, our proposed methods PreDeToR (-$\tau$) do not need to know the action representations or the reward structure and learn these in-context with the help of offline data. The PreDeToR (-$\tau$) predict the reward for the next action of each action during pretraining and can generalize well in-context in several regimes spanning low-data, new actions, and structured bandit settings like linear, non-linear, bilinear, latent bandits. The PreDeToR (-$\tau$) outperforms other in-context algorithms like AD, DPT-greedy in most of the experiments. Finally, we theoretically analyze PreDeToR-$\tau$ and show that pretraining it in $M$ source tasks leads to a low expected excess error on a target task drawn from the same task distribution $\mathcal{T}$. In future, we want to extend our PreDeToR (-$\tau$) to MDP setting (Sutton and Barto, 2018; Agarwal et al., 2019), and constraint MDP setting (Efroni et al., 2020; Gu et al., 2022).
@@ A.1 Related Works
In-context decision making [Laskin et al., 2022, Lee et al., 2023] has emerged as an attractive alternative in Reinforcement Learning (RL) compared to updating the model parameters after collection of new data [Mnih et al., 2013, François-Lavet et al., 2018]. In RL the contextual data takes the form of state-action-reward tuples representing a dataset of interactions with an unknown environment (task). In this paper, we will refer to this as the in-context data. Recall that in many real-world settings, the underlying task can be structured with correlated features, and the reward can be highly non-linear. So specialized bandit algorithms fail to learn in these tasks. To circumvent this issue, a learner can first collect in-context data consisting of just action indices $I_{t}$ and rewards $r_{t}$. Then it can leverage the representation learning capability of deep neural networks to learn a pattern across the in-context data and subsequently derive a near-optimal policy [Lee et al., 2023, Mirchandani et al., 2023]. We refer to this learning framework as an in-context decision-making setting.
@@
The in-context decision-making setting of Sinii et al. [2023] also allows changing the action space by learning an embedding over the action space yet also requires the optimal action during training. In contrast we do not require the optimal action as well as show that we can generalize to new actions without learning an embedding over them. Similarly, Lin et al. [2023] study the in-context decision-making setting of Laskin et al. [2022], Lee et al. [2023], but they also require a greedy approximation of the optimal action. The Ma et al. [2023] also studies a similar setting for hierarchical RL where they stitch together sub-optimal trajectories and predict the next action during test time. Similarly, Liu et al. [2023c] studies the in-context decision-making setting to predict action instead of learning a reward correlation from a short horizon setting. In contrast we do not require a greedy approximation of the optimal action, deal with short horizon setting and changing action sets during training and testing, and predict the estimated means of the actions instead of predicting the optimal action. A survey of the in-context decision-making approaches can be found in Liu et al. [2023a].
@@
In the in-context decision-making setting, the learning model is first trained on supervised input-output examples with the in-context data during training. Then during test time, the model is asked to complete a new input (related to the context provided) without any update to the model parameters [Xie et al., 2021, Min et al., 2022]. Motivated by this, Lee et al. [2023] recently proposed the Decision Pretrained Transformers (DPT) that exhibit the following properties: (1) During supervised pretraining of DPT, predicting optimal actions alone gives rise to near-optimal decision-making algorithms for unforeseen task during test time. Note that DPT does not update model parameters during test time and, therefore, conducts in-context learning on the unforeseen task. (2) DPT improves over the in-context data used to pretrain it by exploiting latent structure. However, DPT either requires the optimal action during training or if it needs to approximate the optimal action. For approximating the optimal action, it requires a large amount of data from the underlying task.
@@
At the same time, learning the underlying data pattern from a few examples during training is becoming more relevant in many domains like chatbot interaction [Madotto et al., 2021, Semnani et al., 2023], recommendation systems, healthcare [Ge et al., 2022, Liu et al., 2023b], etc. This is referred to as few-shot learning. However, most current RL decision-making systems (including in-context learners like DPT) require an enormous amount of data to learn a good policy.
@@
The in-context learning framework is related to the meta-learning framework [Bengio et al., 1990, Schaul and Schmidhuber, 2010]. Broadly, these techniques aim to learn the underlying latent shared structure within the training distribution of tasks, facilitating faster learning of novel tasks during test time. In the context of decision-making and reinforcement learning (RL), there exists a frequent choice regarding the specific ’structure’ to be learned, be it the task dynamics [Fu et al., 2016, Nagabandi et al., 2018, Landolfi et al., 2019], a task context identifier [Rakelly et al., 2019, Zintgraf et al., 2019, Liu et al., 2021], or temporally extended skills and options [Perkins and Precup, 1999, Gupta et al., 2018, Jiang et al., 2022].
@@
However, as we noted in the Section 1, one can do a greedy approximation of the optimal action from the historical data using a weak demonstrator and a neural network policy [Finn et al., 2017, Rothfuss et al., 2018]. Moreover, the in-context framework generally is more agnostic where it learns the policy of the demonstrator [Duan et al., 2016, Wang et al., 2016, Mishra et al., 2017]. Note that both DPT-greedy and PreDeToR are different than algorithmic distillation [Laskin et al., 2022, Lu et al., 2023] as they do not distill an existing RL algorithm. moreover, in contrast to DPT-greedy which is trained to predict the optimal action, the PreDeToR is trained to predict the reward for each of the actions. This enables the PreDeToR (similar to DPT-greedy) to show to potentially emergent online and offline strategies at test time that automatically align with the task structure, resembling posterior sampling.
@@
As we discussed in the Section 1, in decision-making, RL, and imitation learning the transformer models are trained using autoregressive action prediction [Yang et al., 2023]. Similar methods have also been used in Large language models [Vaswani et al., 2017, Roberts et al., 2019]. One of the more notable examples is the Decision Transformers (abbreviated as DT) which utilizes a transformer to autoregressively model sequences of actions from offline experience data, conditioned on the achieved return [Chen et al., 2021, Janner et al., 2021]. This approach has also been shown to be effective for multi-task settings [Lee et al., 2022], and multi-task imitation learning with transformers [Reed et al., 2022, Brohan et al., 2022, Shafiullah et al., 2022]. However, the DT methods are not known to improve upon their in-context data, which is the main thrust of this paper [Brandfonbrener et al., 2022, Yang et al., 2022b].
@@
Our work is also closely related to the offline RL setting. In offline RL, the algorithms can formulate a policy from existing data sets of state, action, reward, and next-state interactions. Recently, the idea of pessimism has also been introduced in an offline setting to address the challenge of distribution shift [Kumar et al., 2020, Yu et al., 2021, Liu et al., 2020, Ghasemipour et al., 2022]. Another approach to solve this issue is policy regularization [Fujimoto et al., 2019, Kumar et al., 2019, Wu et al., 2019, Siegel et al., 2020, Liu et al., 2019], or reuse data for related task [Li et al., 2020, Mitchell et al., 2021], or additional collection of data along with offline data [Pong et al., 2022]. However, all of these approaches still have to take into account the issue of distributional shifts. In contrast PreDeToR and DPT-greedy leverages the decision transformers to avoid these issues. Both of these methods can also be linked to posterior sampling. Such connections between sequence modeling with transformers and posterior sampling have also been made in Chen et al. [2021], Müller et al. [2021], Lee et al. [2023], Yang et al. [2023].
A.2 Experimental Setting Information

### v2

@@ 6 Empirical Study: Importance of Shared Structure and Introducing New Actions
Invariant actions: We denote the set of actions fixed across the different tasks in the pretraining in-context dataset as $\mathcal{A}^{\text{inv}}$. Therefore these action features $\mathbf{x}(a)\in\mathbb{R}^{d}$ for $a\in\mathcal{A}^{\text{inv}}$ are fixed across the different tasks $m$. Note that these invariant actions help the transformer $\mathrm{TF}_{\mathbf{w}}$ to learn the latent structure and the reward correlation across the different tasks. Therefore, as the structure breaks down, PreDeToR starts performing worse than other baselines.
@@
New actions: We also want to test whether PreDeToR (-$\tau$) exploits shared structure when new actions are introduced that are not seen during training time. To this effect, for each task $m\in[M_{\text{pre}}]$ and $m\in[M_{\text{test}}]$ we introduce $A-|\mathcal{A}^{\text{inv}}|$ new actions. That is both for train and test tasks, we introduce new actions. For each of these new actions $a\in[A-|\mathcal{A}^{\text{inv}}|]$ we choose the features $\mathbf{x}(m,a)$ randomly from $\mathcal{X}\subseteq\mathbb{R}^{d}$. Note the transformer now trains on a dataset $\mathcal{H}_{m}\subseteq\mathcal{D}_{\text{pre}}\neq\mathcal{D}_{\text{test}}$.
(a) 00 new action
(b) $1$ new action
(c) $5$ new actions
(d) $10$ new actions
@@
Figure 4: Linear new action experiments. The horizontal axis is the number of rounds. Confidence bars show one standard error.
(a) 00 new action
(b) $1$ new action
(c) $5$ new actions
(d) $10$ new actions
Figure 5: Non-linear new action experiments with non-linear setting.
@@
Finding 4: Shared structure across the tasks is important to learn the reward structure.
@@
Experimental Result: We observe these outcomes in Figure 4 and Figure 5. We consider the linear and non-linear bandit setting of horizon $n=50$, $M_{\text{pre}}=100000$, $M_{\text{test}}=200$, $A=10$, and $d=2$. Here during data collection and during collecting the test data, we randomly select between $0,1,5$, and $10$ new actions from $\mathbb{R}^{d}$ for each task $m$. So the number of invariant actions is $|\mathcal{A}^{\text{inv}}|\in\{10,5,1,0\}$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. From Figure 4(a), 4(b), 4(c), and 4(d), we observe that when the number of invariant actions is less than PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, and AD. Observe that PreDeToR (-$\tau$) matches LinUCB and has lower regret than DPT-greedy, and AD when $\mathcal{A}^{\text{inv}}|\in\{10,5,1\}$. This shows that PreDeToR (-$\tau$) is able to exploit the latent linear structure of the underlying tasks. However, as the number of invariant actions decreases we see that PreDeToR(-$\tau$) performance drops and becomes similar to the unstructured bandits Thomp. We also show in Section A.3 that in $K$-armed bandit setting when there is no structure across arms PreDeToR (-$\tau$) matches the performance of the demonstrator.
Similarly in Figure 5(a), 5(b), 5(c), and 5(d) we show the performance of PreDeToR in the non-linear bandit setting. Observe that LinUCB, MLinGreedy fails to perform well in this non-linear setting due to their assumption of linear rewards. Again note that PreDeToR (-$\tau$) has lower regret than DPT-greedy, and AD when $\mathcal{A}^{\text{inv}}|\in\{10,1\}$. This shows that PreDeToR (-$\tau$) is able to exploit the latent linear structure of the underlying tasks. However, as the number of invariant actions decreases we see that PreDeToR(-$\tau$) performance drops and becomes similar to AD.
7 Data Collection Analysis
In this section, we analyze the performance of PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LinUCB when the weak demonstrator $\pi^{w}$ is Thomp, LinUCB, or Uniform. We again consider the linear bandit setting discussed in Section 4. We show the cumulative regret by the above baselines in Figure 6(a), 6(b), and 6(b) when data is collected through Thomp, LinUCB, and Uniform respectively. We first state the main finding below:
Finding 5: The PreDeToR (-$\tau$) excels in exploiting the underlying latent structure and reward correlation from in-context data when the data diversity is high.
(a) Thomp data collection
(b) LinUCB data collection
(c) Uniform data collection
@@ 7 Data Collection Analysis
We also empirically study the test performance of PreDeToR (-$\tau$) in $K$-armed bandit setting when there is no structure across arms in Section A.3, against the original DPT in Section A.3, in other non-linear bandit settings such as bilinear bandits (Section A.4), latent bandits (Section A.5), draw a connection between PreDeToR and Bayesian estimators (Section A.6), and perform sensitivity and ablation studies in Section A.7, A.9, A.10, A.11. Due to space constraints, we refer the interested reader to the relevant section in the appendices.
8 Theoretical Analysis of Generalization
In this section, we present a theoretical analysis of how PreDeToR-$\tau$ generalizes to an unknown target task given a set of source tasks. We observe that PreDeToR-$\tau$’s performance hinges on a low excess error on the predicted reward of the actions of the unknown target task based on the in-context data. Thus, in our analysis, we show that, in low-data regimes, PreDeToR-$\tau$ has a low expected excess risk for the unknown target task as the number of source tasks increases. This is summarized as follows:
Finding 6: PreDeToR (-$\tau$) has a low expected excess risk for the unknown target task as the number of source tasks increases. Moreover, the transfer learning risk of PreDeToR-$\tau$ (once trained on the $M$ source tasks) scales with $\widetilde{O}(1/\sqrt{M})$.
@@ 8 Theoretical Analysis of Generalization
To show this, we proceed as follows: Suppose we have the training data set $\mathcal{H}_{\mathrm{all}}=\{\mathcal{H}_{m}\}_{m=1}^{M_{\text{pre}}},$ where the task $m\sim\mathcal{T}$ with a distribution $\mathcal{T}$ and the task data $\mathcal{H}_{m}$ is generated from a distribution $\mathcal{D}_{\text{pre}}(\cdot|m).$ For illustration purposes, here we consider the training data distribution $\mathcal{D}_{\text{pre}}(\cdot|m)$ where the actions are sampled following soft-LinUCB (a stochastic variant of LinUCB) (Chu et al., 2011). Given the loss function in Equation 3, we can define the task $m$ training loss of PreDeToR-$\tau$ as ^$\widehat{\mathcal{L}}_{m}(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}})=\frac{1}{n}\sum_{t=1}^{n}\ell(r_{m,t},\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\widehat{r}_{m,t}(I_{m,t})|\mathcal{H}^{t}_{m}))=\frac{1}{n}\sum_{t=1}^{n}(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\widehat{r}_{m,t}(I_{m,t})|\mathcal{H}^{t}_{m})-r_{m,t})^{2}$. We drop the notation $\mathbf{\Theta},\mathbf{r}$ from $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}$ for simplicity and let $M=M_{\text{pre}}$. We define
@@
Note that for in-context learning, a training sample $\left(I_{t},r_{t}\right)$ impacts all future decisions of the algorithm from time step $t+1$ to $n$. Therefore, we need to control the stability of the input perturbation of the learning algorithm learned by the transformer. We introduce the following stability condition.
@@
(Error stability (Bousquet & Elisseeff, 2002; Li et al., 2023)). Let $\mathcal{H}=\left(I_{t},r_{t}\right)_{t=1}^{n}$ be a sequence in $[A]\times[0,1]$ with $n\geq 1$ and $\mathcal{H}^{\prime}$ be the sequence where the $t^{\prime}$th sample of $\mathcal{H}$ is replaced by $\left(I_{t}^{\prime},r_{t}^{\prime}\right)$. Error stability holds for a distribution $(I,r)\sim\mathcal{D}$ if there exists a $K>0$ such that for any $\mathcal{H},\left(I_{t}^{\prime},r_{t}^{\prime}\right)\in([A]\times[0,1]),t\leq n$, and $\mathrm{TF}\in\mathrm{Alg}$, we have
$\big{\lvert}\mathbb{E}_{(I,r)}\left[\ell(r,\mathrm{TF}(\widehat{r}(I)|\mathcal{H}))-\ell\left(r,\mathrm{TF}(\widehat{r}(I)|\mathcal{H}^{\prime})\right)\right]\big{\rvert}\,\leq\tfrac{K}{n}.$
@@
(PreDeToR risk) Suppose error stability 8.1 holds and assume loss function $\ell(\cdot,\cdot)$ is $C$-Lipschitz for all $r_{t}\in[0,B]$ and horizon $n\geq 1$. Let $\widehat{\mathrm{TF}}$ be the empirical solution of (ERM) and $\mathcal{N}(\mathcal{A},\rho,\epsilon)$ be the covering number of the algorithm space $\mathrm{Alg}$ following Definition C.2 and C.3. Then with probability at least $1-2\delta$, the excess MTL risk of PreDeToR-$\tau$ is bounded by
$\mathcal{R}_{\mathrm{MTL}}(\widehat{\mathrm{TF}})\leq 4\tfrac{C}{\sqrt{nM}}+2(B+K\log n)\sqrt{\tfrac{\log(\mathcal{N}(\mathrm{Alg},\rho,\varepsilon)/\delta)}{cnM}},$
where $\mathcal{N}(\mathrm{Alg},\rho,\varepsilon)$ is the covering number of transformer $\widehat{\mathrm{TF}}$ and $\epsilon=1/\sqrt{nM}$.
The proof of Theorem 8.2 is provided in Section C.1. From Theorem 8.2 we see that in low-data regime with a small horizon $n$, as the number of tasks $M$ increases the MTL risk decreases. We further discuss the stability factor $K$ and covering number $\mathcal{N}(\mathrm{Alg},\rho,\varepsilon)$ in Remark C.4, and C.5.
@@ 9 Conclusions, Limitations and Future Works
In this paper, we studied the supervised pretraining of decision transformers in the multi-task structured bandit setting when the knowledge of the optimal action is unavailable. Our proposed methods PreDeToR (-$\tau$) do not need to know the action representations or the reward structure and learn these with the help of offline data. PreDeToR (-$\tau$) predict the reward for the next action of each action during pretraining and can generalize well in-context in several regimes spanning low-data, new actions, and structured bandit settings like linear, non-linear, bilinear, latent bandits. The PreDeToR (-$\tau$) outperforms other in-context algorithms like AD, DPT-greedy in most of the experiments. Finally, we theoretically analyze PreDeToR-$\tau$ and show that pretraining it in $M$ source tasks leads to a low expected excess error on a target task drawn from the same task distribution $\mathcal{T}$. In the future, we want to extend our PreDeToR (-$\tau$) to the MDP setting (Sutton & Barto, 2018; Agarwal et al., 2019), and constrained MDP setting (Efroni et al., 2020; Gu et al., 2022).
@@ A.1 Related Works
In-context decision making (Laskin et al., 2022; Lee et al., 2023) has emerged as an attractive alternative in Reinforcement Learning (RL) compared to updating the model parameters after collection of new data (Mnih et al., 2013; François-Lavet et al., 2018). In RL the contextual data takes the form of state-action-reward tuples representing a dataset of interactions with an unknown environment (task). In this paper, we will refer to this as the in-context data. Recall that in many real-world settings, the underlying task can be structured with correlated features, and the reward can be highly non-linear. So specialized bandit algorithms fail to learn in these tasks. To circumvent this issue, a learner can first collect in-context data consisting of just action indices $I_{t}$ and rewards $r_{t}$. Then it can leverage the representation learning capability of deep neural networks to learn a pattern across the in-context data and subsequently derive a near-optimal policy (Lee et al., 2023; Mirchandani et al., 2023). We refer to this learning framework as an in-context decision-making setting.
@@
The in-context decision-making setting of Sinii et al. (2023) also allows changing the action space by learning an embedding over the action space yet also requires the optimal action during training. In contrast we do not require the optimal action as well as show that we can generalize to new actions without learning an embedding over them. Similarly, Lin et al. (2023) study the in-context decision-making setting of Laskin et al. (2022); Lee et al. (2023), but they also require a greedy approximation of the optimal action. The Ma et al. (2023) also studies a similar setting for hierarchical RL where they stitch together sub-optimal trajectories and predict the next action during test time. Similarly, Liu et al. (2023c) studies the in-context decision-making setting to predict action instead of learning a reward correlation from a short horizon setting. In contrast we do not require a greedy approximation of the optimal action, deal with short horizon setting and changing action sets during training and testing, and predict the estimated means of the actions instead of predicting the optimal action. A survey of the in-context decision-making approaches can be found in Liu et al. (2023a).
@@
In the in-context decision-making setting, the learning model is first trained on supervised input-output examples with the in-context data during training. Then during test time, the model is asked to complete a new input (related to the context provided) without any update to the model parameters (Xie et al., 2021; Min et al., 2022). Motivated by this, Lee et al. (2023) recently proposed the Decision Pretrained Transformers (DPT) that exhibit the following properties: (1) During supervised pretraining of DPT, predicting optimal actions alone gives rise to near-optimal decision-making algorithms for unforeseen task during test time. Note that DPT does not update model parameters during test time and, therefore, conducts in-context learning on the unforeseen task. (2) DPT improves over the in-context data used to pretrain it by exploiting latent structure. However, DPT either requires the optimal action during training or if it needs to approximate the optimal action. For approximating the optimal action, it requires a large amount of data from the underlying task.
@@
At the same time, learning the underlying data pattern from a few examples during training is becoming more relevant in many domains like chatbot interaction (Madotto et al., 2021; Semnani et al., 2023), recommendation systems, healthcare (Ge et al., 2022; Liu et al., 2023b), etc. This is referred to as few-shot learning. However, most current RL decision-making systems (including in-context learners like DPT) require an enormous amount of data to learn a good policy.
@@
The in-context learning framework is related to the meta-learning framework (Bengio et al., 1990; Schaul & Schmidhuber, 2010). Broadly, these techniques aim to learn the underlying latent shared structure within the training distribution of tasks, facilitating faster learning of novel tasks during test time. In the context of decision-making and reinforcement learning (RL), there exists a frequent choice regarding the specific ’structure’ to be learned, be it the task dynamics (Fu et al., 2016; Nagabandi et al., 2018; Landolfi et al., 2019), a task context identifier (Rakelly et al., 2019; Zintgraf et al., 2019; Liu et al., 2021), or temporally extended skills and options (Perkins & Precup, 1999; Gupta et al., 2018; Jiang et al., 2022).
@@
However, as we noted in the Section 1, one can do a greedy approximation of the optimal action from the historical data using a weak demonstrator and a neural network policy (Finn et al., 2017; Rothfuss et al., 2018). Moreover, the in-context framework generally is more agnostic where it learns the policy of the demonstrator (Duan et al., 2016; Wang et al., 2016; Mishra et al., 2017). Note that both DPT-greedy and PreDeToR are different than algorithmic distillation (Laskin et al., 2022; Lu et al., 2023) as they do not distill an existing RL algorithm. moreover, in contrast to DPT-greedy which is trained to predict the optimal action, the PreDeToR is trained to predict the reward for each of the actions. This enables the PreDeToR (similar to DPT-greedy) to show to potentially emergent online and offline strategies at test time that automatically align with the task structure, resembling posterior sampling.
@@
As we discussed in the Section 1, in decision-making, RL, and imitation learning the transformer models are trained using autoregressive action prediction (Yang et al., 2023). Similar methods have also been used in Large language models (Vaswani et al., 2017; Roberts et al., 2019). One of the more notable examples is the Decision Transformers (abbreviated as DT) which utilizes a transformer to autoregressively model sequences of actions from offline experience data, conditioned on the achieved return (Chen et al., 2021; Janner et al., 2021). This approach has also been shown to be effective for multi-task settings (Lee et al., 2022), and multi-task imitation learning with transformers (Reed et al., 2022; Brohan et al., 2022; Shafiullah et al., 2022). However, the DT methods are not known to improve upon their in-context data, which is the main thrust of this paper (Brandfonbrener et al., 2022; Yang et al., 2022b).
@@
Our work is also closely related to the offline RL setting. In offline RL, the algorithms can formulate a policy from existing data sets of state, action, reward, and next-state interactions. Recently, the idea of pessimism has also been introduced in an offline setting to address the challenge of distribution shift (Kumar et al., 2020; Yu et al., 2021; Liu et al., 2020; Ghasemipour et al., 2022). Another approach to solve this issue is policy regularization (Fujimoto et al., 2019; Kumar et al., 2019; Wu et al., 2019; Siegel et al., 2020; Liu et al., 2019), or reuse data for related task (Li et al., 2020; Mitchell et al., 2021), or additional collection of data along with offline data (Pong et al., 2022). However, all of these approaches still have to take into account the issue of distributional shifts. In contrast PreDeToR and DPT-greedy leverages the decision transformers to avoid these issues. Both of these methods can also be linked to posterior sampling. Such connections between sequence modeling with transformers and posterior sampling have also been made in Chen et al. (2021); Müller et al. (2021); Lee et al. (2023); Yang et al. (2023).
A.2 Experimental Setting Information and Details of Baselines
In this section, we describe in detail the experimental settings and some baselines.
A.2.1 Experimental Details